A puck with a mass of 0.170kg has velocity of +90km/H. The momentum of the puck is ??


1
Expert's answer
2021-05-18T17:29:48-0400
"p=mv,""p=0.17\ kg\cdot90\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s}=4.25\ \dfrac{kg\cdot m}{s}."